OpenStudy (anonymous):

The relationship between the fixed focal length F of a camera, the distance u from the object being photographed to the lens, and the distance v from the lens to the film is given by the equation. If the focal length is 80 millimeters, express v as a function of u. (1/F)=(1/u)+(1/v)

OpenStudy (across):

\[\frac{1}{F}=\frac{1}{u}+\frac{1}{v},\]\[\frac{1}{80}=\frac{1}{u}+\frac{1}{v},\]\[\frac{1}{80}-\frac{1}{u}=\frac{1}{v},\]\[\frac{u-80}{80u}=\frac{1}{v},\]\[v=\frac{80u}{u-80}.\]